Can you please help me with theses 2 questions?

1). Find three arithmetic means between 11 and 31. (Enter your answers as a comma-separated list.)

2). Evaluate the piecewise-defined function.

f(x) =

3 if x < 0
3 − x     if 0 ≤ x < 3
x + 3     if x ≥ 3

(a) f(4/3)=

(b) f(3)=
Answer #1

0 Three arithmetic means between 11 and 31. 11,2,,X2,83, 31 ä, = first term = 11 95=3) an = 9, + (n-1)d 95= 917(5-1)d 3) = 11fox) 3 {: if xco 3-2 if 054<3 if xz3 xt3 @ f() where you the value of a x = 42 4 satirofying the condution Osacz so function